OpenStudy (naga447):

I will medal the first person to help!!! Jack invested some money in a bank at a fixed rate of interest compounded annually. The equation below shows the value of his investment after x years: f(x) = 300(1.02)x What was the average rate of change of the value of Jack's investment from the third year to the fifth year? 6.43 dollars per year 8.24 dollars per year 12.86 dollars per year 14.26 dollars per year

OpenStudy (naga447):

Okay how would I find the difference because the y values for year 3 is 318.36 and year 5 is 331.22

OpenStudy (timmyk14):

Alright, now I see. You need to solve this function for the following years. The 3rd year, the 4th year, and the 5th year. Next you want to add the values and then divide the sum by 3. That's how you find the average of a set or numbers. Still need help?

OpenStudy (timmyk14):

Ok...So this is what you should be doing: F(3) = [300(1.02*3)] F(4) = [300(1.02*4)] F(5) = [300(1.02*5)] Find the values of x in all of those functions and then tell them to me. If you need help finding the values of x in these functions let me know.

OpenStudy (naga447):

f(3) = 918 f(4) = 1224 f(5) =1530

OpenStudy (timmyk14):

Okay, this question is confusing me in many ways. Even after finding the average rate of change the answer isn't one of the choices you provided. I've re-read this question multiple times to see what I might be missing but I can't find anything. I will keep looking, but for now I'm stuck. Sorry, I though I was on the right track, but it turns out I wasn't. Maybe we can get a mod to explain to us what to do..

OpenStudy (timmyk14):

Hey @cwrw238 do you mind helping us with this question. I'm unsure of what to do, and even though this isn't my question I would love to see what I was doing wrong. Thanks!

OpenStudy (timmyk14):

Well until he has a chance to view your question I want to keep trying. So, it says that the function f(x) = 300(1.02)x is the total value of his investment after x years. The question wants us to find the average rate of change of his investment from the 3rd to 5th year. Now we just need to figure out exactly what we need to do, and how we will do it...Let me think for a second

OpenStudy (timmyk14):

Something tells me that we might need to make a graph to find the RATE OF CHANGE...so let me draw one real quick and explain further

OpenStudy (timmyk14):



OpenStudy (timmyk14):

Rate of change is basically the slope of a data set right? If so then all we need to do is find the slope. Remember this though, I'm not completely sure of this...

OpenStudy (naga447):

The Ave. rate of change is 1224 right?

OpenStudy (timmyk14):

No, I was wrong when it came to that :/

OpenStudy (timmyk14):

So we need to do this rise over run What is the "rise over run" of the first set of points? (3,918) and (4,1224)

OpenStudy (naga447):

oh, the equation is y1 - y2 over x2 - x1 right?

OpenStudy (timmyk14):

@cwrw238 We seriously need help. Please.

OpenStudy (naga447):

I have no idea what's happening...

OpenStudy (timmyk14):

I'm researching some stuff. I haven't gotten much sleep lately so I'm really tired and not thinking properly.

OpenStudy (naga447):

Okay, but if it's becoming too much of a problem don't worry about it and get some sleep :)

OpenStudy (timmyk14):

Okay, this video might help us with this problem. https://www.youtube.com/watch?v=jqIoqECyY0A Let me know when you finish watching it and we can try and solve this problem.

OpenStudy (timmyk14):

This question is kicking my a** right now. I've tried absolutely everything. I've used pages upon pages in my notebook writing down equations and solving them and it seems as though none of them are right. I've done what the video said to do and I found the rate of change which is 306. I already knew that because 918+306=1224, and 1224+306=1530. I'm unsure of what this question is even asking and I feel terrible that I can't help you. I've done all I could, ask some of the Math Professors and moderators for help. They might know what this question is asking and might be able to help you better than what I've done. I've searched for videos on the subject and I've gone to yahoo to see what I'm supposed to do and in the end none of them have worked. Sorry again, and I hope you find the correct answer, when you do tag me so I can see what you needed to do,

OpenStudy (naga447):

Thank you for your help! but I've finished the home work I'm going to ask my teacher to explain to me what I'm missing, Again thanks for your help!

OpenStudy (timmyk14):

@naga4447 Well, please let me know how your teacher solved this so I can learn from it. Learning is key right xD I love math, and not being able to answer this question seriously irritates me. So yea, I would really appreciate it if you let me know how to solve this :)

OpenStudy (timmyk14):

Whoops, meant to say @naga447